Can anyone solve and explain this?
Write a quadratic equation having the given numbers as solutions:
4i, -4i

Answer by stanbon(75887)

Write a quadratic equation having the given numbers as solutions:
4i, -4i
-------------------
Factor Theorem: If "a" is a root of an equation, x-a is a factor.
-------------------------
You are given two roots:
The corresponding factors are (x-4i) and (x-(-4i)) or (x+4i)
-----
Equation:
f(x) = (x+4i)(x-4i)
f(x) = x^2 +4ix - 4ix - (4i)^2
---
f(x) = x^2 - 16i^2
f(x) = x^2 - 16*(-1)
f(x) = x^2 + 16
=====================
Cheers,
Stan H.
=====================